OpenStudy (anonymous):

The tangent at a point P on the curve x^4 y^5 = a^9 meet the x and y axes in A and B respectively Then AP:PB=?

OpenStudy (anonymous):

x^4 y^4 =a^9 d(x^4 y^4)/dx =d(a^9)/dx d(x^4)/dx *y^4 + x^4 d(y^4)/dx =0 4x^3y^4 +x^4 4y^3 dy/dx=0 dy/dx =-4x^3y^4/(4x^4y^3) dy/dx =-y/x

OpenStudy (anonymous):

Oh...No......There is a mistake in the question Sorry

OpenStudy (anonymous):

It is x^4 y^5 = a^9

OpenStudy (anonymous):

dy/dx = -4y/5x

OpenStudy (anonymous):

x^4 y^5 =a^9 d(x^4 y^5)/dx =d(a^9)/dx d(x^4)/dx *y^5 + x^4 d(y^5)/dx =0 4x^3y^5+x^4 5y^4 dy/dx=0 dy/dx =-4x^3y^5/(5x^4y^4) dy/dx =-4y/5x

OpenStudy (anonymous):

Now, coordinate of P is